Let f: R2 to R defined by f(x,y) =x2+y2 .show that f is differentiable at (x0,y0) element. Of R2 and find gradient of f at (x0,y0).


1
Expert's answer
2021-03-22T08:40:22-0400

Given f: R2 "\to" R

such that f(x,y) = x2 + y2

We have to show that f is differentiable at (x0,y0)

given (x0,y0) be an arbitrary point in R2 . Since f(x0,y0) = x02 + y02

fx(x0,y0) = 2x0 & fy(x0,y0) = 2y0

we find that the linear approximation to f at (x0,y0) is given by :

L(x, y) = x02 + y02 + 2x0(x − x0) + 2y0(y − y0) = 2x0x + 2y0y − x02 −y02 .

Therefore,

f(x, y) − L(x, y) = x2 + y2 − ( 2x0x + 2y0y − x02 −y02 )

= x2 + y2 − 2x0x - 2y0y + x02 + y02 )

= (x − x0)2 + (y − y0)2

It follows that,

"\lim_{(x,y) \to (x_0,y_0)}" "\frac{f(x, y) − L(x, y)
}{ \sqrt{(x − x_0)^2 + (y − y_0)^2}}" = "\lim_{(x,y) \to (x_0,y_0)}" "\frac{(x − x_0)^2 + (y − y_0)^2
}{ \sqrt{(x − x_0)^2 + (y − y_0)^2}}"


= "\lim_{(x,y) \to (x_0,y_0)}""\sqrt{(x − x_0)^2 + (y − y_0)^2}"

= 0


Since the final function to the right of the limit symbol is continuous at (x0,y0). Since the necessary limit is zero, f is differentiable at (x0,y0).


Now to find Gradient of f at (x0,y0)

Given f(x,y) = x2 + y2

( we know that grad(f) = "\nabla f" = < fx , fy > = fx i + fy j )

therefore for given f(x,y),

fx = 2x & fy = 2y

"\nabla f" = < 2x, 2y > = 2x i + 2y j


"\nabla f" at (x0,y0) = < 2x0, 2y0 > = 2x0 i + 2y0 j